Calculate the percent ionization of carbonic acid (H2CO3) in solutions of each of the following concentrations (Ka = 4.3e-07.)



(a) 0.193 M





(b) 0.371 M




(c) 0.783 M
Answer #1

= ut tuco, initial: at equili: H₂CO3 0.193 0.193-2 x ka = 4.3 x 10-7 = 42 0193-X O: 362 +4.3x10-734-0.8299xlot 2= 0.0002878 m/. ionization: 000399 X100 - 0.1075.1. .391 - 4 + 4CO3 - HCO3 initial: 783 at equili: •783- K 2 Ka = 4.3 4/0-7 = 23 •783-3 0